OpenStudy (anonymous):

Please help Joy and Cullen put some money into their money boxes every week. The amount of money (y), in dollars, in their money boxes after a certain amount of time (x), in weeks, is shown by the equations below: Joy: y = 30x + 20 Cullen: y = 20x + 90 After how many weeks will Joy and Cullen have the same amount of money in their money boxes, and what is that amount? 6 weeks, $150 230 weeks, $7 7 weeks, $150 7 weeks, $230

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

y = 30x + 20 20x + 90 = 30x + 20 ... Replace y with 20x+90 20x + 90 - 90 = 30x + 20 - 90 ... Subtract 90 from both sides. 20x + 0 = 30x - 70 20x = 30x - 70 20x = -70 + 30x 20x - 30x = -70 + 30x - 30x ... Subtract 30x from both sides. -10x = -70 + 0x -10x = -70 x = ???

OpenStudy (anonymous):

7 ? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

x = 7, yep

jimthompson5910 (jim_thompson5910):

now plug that value of x into either original equation to find y

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

I recommend you do both equations to see that it's actually the solution

OpenStudy (anonymous):

y = 230 for both @jim_thompson5910

OpenStudy (anonymous):

Oh I understand it now it would be D. 230 dollars and 7 weeks, Thank you.

jimthompson5910 (jim_thompson5910):

x = 7 and y = 230, correct